# A new indeterminacy-based quantum theory

### v4

@@ 2 Main features: four conditions and two hypotheses
The criteria and some others will allow EnDQT to provide the great benefit of giving criteria for when determinate values arise, but without necessarily modifying the fundamental equations of QT like spontaneous collapse theories or adopting a relationalist view or hidden variables. As we can see, systems with the DC and determinate values propagate through interactions. I will call a chain of interactions between systems that propagate the DC, a stable determination chain (SDC). It is stable because it can be seen as a stable process that gives rise to the spread of determinacy between systems.
Note that, as we can anticipate, in order for a system $E$ to decohere another $S$, it has to be interacting with other systems that have the DC. Those systems will have to interact with other systems with the DC, and so on. So, we will need many systems for a system $S$ to be fundamentally decohered by $E$, giving rise to $S$ having a determinate value. These systems will have to interact with $S$, as well as other members of the SDC. This process starts to resemble a process represented by a pragmatic irreversible decoherence model because we have here a great number of systems whose states will be very hard to control.
@@
Returning to our conditions, given an SDC, we run the risk of an infinite regress because it’s unclear where it starts. To circumvent these issues,
@@
CDC3) I will consider that there are two kinds of systems that constitute an SDC,
@@
-Initiator systems or initiators, which are systems that have the DC concerning any system by default (i.e., they have the DC-X for all systems X), i.e., independently of their interactions with other systems. Because of this, initiators are the systems that start SDCs.
@@
-Non-initiator systems are systems that don’t have the DC by default but have it due to their interactions with other systems that have the DC.
@@
So, the (fundamental) decoherence of some system $S$ by an initiator is necessary and sufficient to allow that later system to have a determinate value of some observable $O$ of $S$. Also, $S$ can acquire the DC concerning some other system $S^{\prime}$ if it interacts with this system while it interacts with the initiator. We will see further below a system that is a plausible candidate to be an initiator, and which is widely accepted in cosmology. I will argue that although initiators arose to address the measurement problem, they have the advantage of potentially addressing other problems in the foundations of physics, which is a good sign.
@@
SDCs are represented by directed graphs, which represent the propagation of the DCs, or its potential propagation, which gives rise to systems that belong to it having determinate values. I will represent this interaction between $X$ that leads $Y$ to have a determinate value (together with $X$) and potentially leads $Y$ to have the DC concerning some other system as $X\rightarrow Y$. When $Y$ has the DC-Z, which leads some other system $Z$ to have a determinate value, I will represent it as $X\rightarrow Y\rightarrow Z$ [^23] For simplicity, here I will mostly not care about the distinction between a token network, which represents concrete interactions between systems, and type networks, which represent interactions between types of systems that exist in specific regions of spacetime. In some DAGs that aim to depict the whole situation, the systems with only directed arrows towards them represent systems that have the DC but won’t end up transmitting it to other systems. An SDC ends when it reaches these systems. The nodes with no directed arrows towards them represent the initiators.
Let’s consider a simple and idealized example where, once again, we can neglect the intrinsic evolution of systems due to their strong interaction. This example will involve systems $A$, $B$, and $C$, where $A$ is an initiator, in a toy mini-universe where the SDC that will be formed has the following structure, $A\rightarrow B\rightarrow C$. Let’s assume that $C$ interacts with $B$ while $B$ is interacting with $A$, where the interactions between $A$ and $B$ started first so that $B$ can end up transmitting the DC to $C$.
However, let’s suppose that $B$ and $C$ start interacting while $A$ and $B$ interact. However, when $B$ and $C$ begin interacting, let’s assume that we can neglect the evolution of the quantum states of $B$ while $A$ and $B$ interact, such that we can idealize that $B$ and $C$ start interacting only when the interaction between $A$ and $B$ ends. Thus, we can just analyze the evolution of the quantum states of $A$ while $A$ and $B$ are interacting, where this interaction ends approximately at $t^{\prime}.$ [^24] Whether this assumption is a requirement for SDCs to develop or just a simplification needs to be further studied. We have then the following interaction between $A$ and $B$,
@@
If $\left\langle E_{0}\left(t^{\prime}\right)\mid E_{1}\left(t^{\prime}\right)\right\rangle_{\text{A SDC}}\approx 0$ and $\left\langle E_{1}\left(t^{\prime}\right)\mid E_{0}\left(t^{\prime}\right)\right\rangle_{\text{A SDC}}\approx 0$ quasi-irreversibly when $A$ and $B$ end their interaction, we infer that $B$ has a determinate value of the observable $monitored$ by $A$ at $t^{\prime}$ that arises from their interaction (i.e., let’s assume that is either $E^{\prime}_{0}$ or $E^{\prime}_{1}$), and acquires the DC. Let’s assume that $B$ has a determinate value $E^{\prime}_{0}$. Now, let’s consider the interaction between $B$ and $C$, which (given our idealization) starts when the interaction between $A$ and $B$ ends. Let’s assume that, given their interaction Hamiltonian, it ends at $t^{\prime\prime}$,
@@
The evolution of the interaction between $B$ and $C$ could be analyzed via the reduced density operator $\rho_{C}(t)$. Their interaction will allow $C$ to have a determinate value ($\uparrow$ or $\downarrow$) at $t^{\prime\prime}$ and the DC if $\langle E_{0}^{\prime\uparrow}\left(t^{\prime\prime}\right)|E_{0}^{\downarrow}\left(t^{\prime\prime}\right)\rangle_{B}\approx 0$ and $\langle E_{0}^{\prime\downarrow}\left(t^{\prime\prime}\right)|E_{0}^{\uparrow}\left(t^{\prime\prime}\right)\rangle_{B}\approx 0$ quasi-irreversibly when $B$ and $C$ end their interaction. $B$ will have a determinate value at $t^{\prime\prime}$ that arises from its interaction with $C$ where the possible values that it can have are represented via the eigenvalues of the observable that $|E_{0}^{\prime\uparrow}\left(t^{\prime\prime}\right)\rangle_{B}$ and $|E_{0}^{\prime\downarrow}\left(t^{\prime\prime}\right)\rangle_{B}$ are eigenstates of. Furthermore, $C$ could transmit the DC to another system of that universe if it interacted with it before the interaction with $B$ ends. Note that since system $A$ is an initiator, its ability to give rise to other systems having determinate values and provide the DC doesn’t depend on the interactions with other systems.
In decoherence models, the environment of a system is typically composed of many subsystems. So, it’s more realistic and plausible to assume that
For instance, let’s consider that instead of $S_{2}$ above, we have a subsystem $S_{2}^{i}$ of $S_{2}$ for some $i$ where $S_{2}^{i}$ is not able to decohere $S_{3}$ alone, but $S_{2}$ is. $S_{2}$ would just be able to give rise to $S_{3}$ having a determinate value, having the DC, if its subsystems $S_{2}^{i}$ for all $i$ interacted with subsystems of $S_{1}$, acquiring the DC, where $S_{1}$ and its subsystems have the DC. So, subsystems of a system, such as $S_{2}^{i}$ for all $i$ are spacelike separated from each other and are considered to form a "cause" for the "common effect," which is a system $S_{3}$ having a determinate value of one of its observables. Each subsystem of $S_{2}$ would also have another determinate value when $S_{3}$ has a determinate value. These interactions are represented by a directed graph with "colliders" (Figure 1). We can also simplify the structure of the above graph by treating $S_{2}$ as a whole, neglecting its subsystems (Figure 2).
CDC4) puts further constraints on the structure and persistence of SDCs, whose elements already have to obey the other CDCs. It’s plausible to consider that typically a system will interact with many systems in spatiotemporal regions, which form a larger system $E$. In order for a system $E$ with its subsystems to decohere another $S$, its subsystems will have to be interacting with other systems that have the DC, and those systems will have to be interacting with other systems, and so on. So, we will need many systems for a system interacting with a larger system $E$ to have a determinate a determinate value due to $E$. Also, the more macroscopic $S$ is, the more subsystems in principle has, and so the more systems belonging to SDCs need to be interacting with $S$ in order for it to have a determinate value and the DC. Given the number of systems that seem to be involved, this process resembles, even more, a process represented by a pragmatic irreversible decoherence model because we will have many systems involved in the interaction that leads $S$ to have a determinate value, which will be hard to control, not to mention that SDCs give rise to indeterministic processes.
@@
Figure 1: Directed graph that involves a common effect (i.e., a "collider") that represents the transmission of the DC between systems.
@@
Figure 2: Directed graph involving common effects like the previous one, but that includes the whole system $S_{2}$.
As we can begin to see more clearly, EnDQT provides the benefit of providing a no-hidden-variable criterion for systems to have determinate values without modifying the fundamental equations of QT or appealing to relationalism. CDC1)-CDC4) complete my presentation of the conservative determination conditions.
Now, given the CSCs we can see how EnDQT justifies the assumptions $i^{*}-iii^{*})$ of decoherence models. Regarding $i^{*})$, it’s important that the Hamiltonian of interaction assumes a specific form to account for decoherence (given by the above commutativity criterion), which depends on the observable of the environment and the system because we can regard it as representing the law-like dynamics of SDCs, which leads to determinate values. The observables of the environment that appear in the Hamiltonian are also constrained by the interactions of the environmental systems of the environment with members of SDCs. Regarding $ii^{*})$, the initial states of the systems (and associated coupling constants) are randomly distributed (which gives rise to the overlap terms going quasi-irreversibly to zero in decoherence models) because such initial environmental systems also need to belong to an SDC that indeterministically gives rise to determinate values and, hence, to a random distribution of their quantum states (and associated coupling constants). [^25] What about the quantum states of the initiators? For example, the initiators can also interact among themselves and give rise to these randomly distributed quantum states or that can just be a brute fact. More on initiators below. Finally, regarding $iii^{*})$, what explains why and how systems (composed of many subsystems) that have the DC give rise to determinate values in this law-like way, where this is represented via fundamental decoherence, can be elegantly (as we will see) traced back to the initiators (which have the DC by default) led them to have to DC and to give rise to systems having determinate values in this law-like way (more on this below). This is one more benefit of adopting EnDQT with its CDCs: it addresses some ad-hoc or strange features of decoherence by seeing it as a tool to represent the dynamics of SDCs.
@@
I have been assuming that we can rely on fundamental decoherence to give an account and infer how determinate values arise via SDCs. I also explained that it seems to represent phenomena that resemble the one represented by pragmatic irreversible decoherence models. Remember that pragmatic irreversible decoherence models are models that involve situations where it’s impossible to reverse the process represented by them because they concern open/inaccessible environments or environments with many systems. As it should be clearer, the models of decoherence that represent the interactions between systems, starting with the initiators are the fundamental decoherence models. The example above and CDC1)-CDC4) involve these models. More concretely, roughly, fundamental decoherence represents interactions that, in a certain situation involving an environment with systems with a DC, given an appropriate and local Hamiltonian, render systems with determinate values. Importantly, in order for a system $S$ to have a DC and lead other systems to have determinate values, it’s plausible that it will typically have to belong to an environment with many systems that propagate the DC, which seems to lead to a process that is hard to reverse/control due to these many systems that have determinate values indeterministically. Furthermore, the more macroscopic is $S$, there harder this seems to be. So, it seems that we can use pragmatic irreversible decoherence models to infer when the processes represented by fundamental decoherence occur.
However, as we will see, more needs to be said about the behavior of SDCs and fundamental decoherence to ground the typically used pragmatic irreversible decoherence models as a proper inferential tool for inferring when systems have determinate values in situations where we have not followed the interactions between systems since the beginning of an SDC.
@@
As a reminder, I have called the models of decoherence that don’t necessarily track the interactions involving systems with the DC pragmatic decoherence models. Another kind of pragmatic decoherence model is what I will call the pragmatic reversible decoherence models. These are models that represent a process that apparently involves decoherence (in the sense that the overlap terms go quasi-irreversibly to zero), but it’s considered that someone in some privileged position could reverse this process via operations on the system and environment (what is called recoherence). This reversible process often occurs inside isolated environments or situations where the environmental degrees of freedom don’t become inaccessible to be reversed due to their in-practice isolation. These processes represented by these models aren’t what we typically consider to be decoherence.
@@
The distinction between what constitutes a pragmatic reversible and irreversible model may be ambiguous if we are not careful. The Wigner’s friend scenario [^26] Wigner (1995). is a situation where this ambiguity becomes manifest. Suppose a large isolated lab (as big as we want) with a human agent inside (a "friend"). The lab is isolated in such a way that Wigner outside the lab can unitarily manipulate the state of the friend plus their target system in a superposition that the friend measures, treating both as being in an entangled state. So, in this case, we could have a lab as large as we want and many systems getting entangled with the friend and the target system for a long time. However, we wouldn’t consider that there was decoherence of the system by the friend if we were Wigner. Wigner could still unitarily reverse the state of the friend and their system. He would rather consider treating the friend plus target system interaction via a reversible decoherence pragmatic model. However, if the lab were open, he would treat their interaction via an irreversible decoherence pragmatic model. So, how do we exactly distinguish a reversible decoherence process from an irreversible one, since the reversible could also involve many systems? Let’s call this problem the Wigner’s friend ambiguity problem. As the discussion in the previous paragraphs suggests, if we want to address the measurement problem without changing the fundamental equations of QT, adding hidden variables to it, or adopting a relationalist view (more on this below), it seems that a way to deal with this problem is via paying attention to what constitutes an open environment and how it relates with SDCs.
@@
I have pointed out how fundamental decoherence involving members of an SDC seems to lead to processes represented by the pragmatic irreversible decoherence models. The first hypothesis aims to deal with the above ambiguity problem and ground the success of the pragmatic irreversible decoherence models in helping account for determinate values arising via SDCs in open environment situations, as opposed to the reversible models. Thus, I will pose that
@@
The SDCs in our world are widespread in such a way that pragmatic irreversible decoherence models in open environments track the interactions between systems that belong to SDCs, but there can also exist processes represented via reversible decoherence pragmatic models, where the latter are tracking the interactions between systems that don’t belong to SDCs (SDCs-decoherence hypothesis).
@@
In other words, what the SDCs-decoherence hypothesis says is that the SDCs in our world are such that the open environments involved in irreversible decoherence pragmatic models will give rise to many interactions between the target system and systems that have the DC, in such a way that we can reliably approximate fundamental decoherence models by these pragmatic irreversible decoherence models. From now on, I will just assume that processes represented by pragmatic irreversible decoherence models occur in open environments.
@@
On the other hand, it also hypothesizes that SDCs in our world are such that in isolated/controlled situations, we might be able to isolate the target system from interacting with members of SDCs, only interacting with systems that don’t have the DC. It all depends on the history of the SDCs that could interact with the target system in those situations and on our ingenuity in shielding the target system from interacting with SDCs (more on this below).

### v5

@@ 2.1 The four conditions
I also want to emphasize that these criteria and a few others will allow EnDQT to provide the great benefit of giving conditions for when determinate values arise, but without modifying the fundamental equations of QT like spontaneous collapse theories or adopting a relationalist view or hidden variables. As we can see, systems with the DC and determinate values propagate through interactions. I will call a chain of interactions between systems that propagate the DC or just give rise to systems having determinate values a stable determination chain (SDC). It is called SDC because the process that gives rise to systems having determinate values can be seen as a process that, in order to occur, needs to be stable in the sense that it stably obeys CDC1*). More concretely, in order to infer that systems have determinate values of some observable, it is necessary that the overlap terms of the quantum states of the environment go $stably$ to zero. Also, given CDC1*) again, it is also necessary that systems that interact with systems that are going over this process, are $stably$ not disturbed from going over this process.
One may object that contrary to what is claimed, EnDQT is really modifying the equations because it’s posing a time where an indeterministic process occurs and stops a process represented deterministically and unitarily. Note, however, that the point is that we won’t have to modify the fundamental equations of QT to infer this process contrary to, for example, as I have said, spontaneous collapse theories.
@@
Returning to our conditions, given an SDC, we run the risk of an infinite regress because it is unclear where and when it starts. To circumvent these issues, I will consider that
@@
CDC3) There are two kinds of systems that constitute an SDC, or two interrelated kinds of roles they have
@@
-Initiator systems or initiators, which are systems that a) have the DC concerning any system by default (i.e., they always have the DC-X for any system X), i.e., independently of their interactions with other systems. Or b) they are the first systems that have the DC or the ones that we assign the DC initially in our models. Because of this, initiators are the systems that start SDCs.
@@
-Non-initiator systems are systems that don’t have the DC concerning a system by default but have it due to their interactions with other systems that have the DC.
@@
So, the (fundamental) decoherence of some system $S$ by an initiator is necessary and sufficient to allow that later system to have a determinate value of some observable $O$ of $S$. Also, $S$ can acquire the DC concerning some other system $S^{\prime}$ if it interacts with this system while it interacts with the initiator. We will see further below a system that is a plausible candidate to be an initiator, and which is widely accepted in cosmology. I will argue that although initiators arose to address the measurement problem, they have the advantage of potentially addressing other problems in the foundations of physics, which is a good sign. Note that initiators of the kind b) are identical to non-initiators. They just end up being the first systems with the DC in the universe. Or if we conceive that the universe existed forever, as well as the DC, they are the systems that we assign the DC to in our models at first. On the other hand, initiators of the kind a) are special systems because they always have the DC concerning any system. It is less parsimonious to postulate initiators of the kind a) than b).
@@
SDCs are represented by directed graphs, which represent the propagation of the DCs, or its potential propagation, which gives rise to systems that belong to it having determinate values. I will represent this interaction between $X$ that leads $Y$ to have a determinate value (together with $X$) and potentially leads $Y$ to have the DC concerning some other system as $X\rightarrow Y$. When $Y$ has the DC-Z, which leads some other system $Z$ to have a determinate value, I will represent it as $X\rightarrow Y\rightarrow Z$ [^34] For simplicity, here I will mostly not care about the distinction between a token network, which represents concrete interactions between systems, and type networks, which represent interactions between types of systems that exist in specific regions of spacetime. In some graphs that aim to depict the whole situation, the systems with only directed arrows towards them represent systems that have the DC but won’t end up transmitting it to other systems. An SDC ends when it reaches these systems. The nodes with no directed arrows towards them represent the initiators.
Let’s turn to the last CDC). In decoherence models, the environment of a system is typically composed of many subsystems. So, it is more realistic and plausible to assume that
For instance, let’s consider that instead of $S_{2}$ above, we have a subsystem $S_{2}^{i}$ of $S_{2}$ for some $i$ where $S_{2}^{i}$ is not able to decohere $S_{3}$ alone, but $S_{2}$ is. $S_{2}$ would just be able to give rise to $S_{3}$ having a determinate value, having the DC, if its subsystems $S_{2}^{i}$ for all $i$ interacted with subsystems of $S_{1}$, acquiring the DC, where $S_{1}$ and its subsystems have the DC. So, subsystems of a system, such as $S_{2}^{i}$ for all $i$ are considered to form a “cause” for the “common effect,” which is a system $S_{3}$ having a determinate value of one of its observables. Each subsystem of $S_{2}$ would also have another determinate value when $S_{3}$ has a determinate value. These interactions are represented by a directed graph with “colliders” (i.e., a series of arrows that point to a single system, see Figure 1). We can also simplify the structure of the above graph by treating $S_{2}$ as a whole, neglecting its subsystems (Figure 2).
CDC4) further constrains the structure and persistence of SDCs, whose elements already have to obey the other CDCs. It is plausible to consider that typically, a system will interact with many systems in spatiotemporal regions, which form a larger system $E$. In order for a system $E$ with its subsystems to decohere another $S$, its subsystems will have to be interacting with other systems that have the DC concerning them, and those latter systems will have to be interacting with other systems, and so on. So, we will need many systems for a system interacting with a larger system $E$ to have a determinate value due to $E$. Also, the more macroscopic $S$ is, the more subsystems in principle has, and so more systems belonging to SDCs need to be interacting with $S$ in order for it to have a determinate value and the DC concerning other systems. This process resembles a process represented by a pragmatic irreversible decoherence model because we will have many systems involved in the interaction that leads $S$ to have a determinate value and perhaps the DC, and these systems will be hard to control. Also, the SDCs give rise to indeterministic processes, which make these interactions impossible to reverse unitarily. [^35] Note that it’s plausible to consider that systems that belong to an SDC having determinate values due to a target system $A$ will depend on which other systems are interacting with $A$ at that time (all of them having a determinate value, constituting single events). Via their influence on $A$ having a determinate value, they also influence each other to have determinate values. Because of this, decoherence of a target system due to a collective of systems allows us to infer the process that gives rise to a collective of systems having determinate values (such as in the case of central spin models, i.e., a model with a system with a spin interacting with multiple systems during a period, see [29] and Appendix A).
CDC1)-CDC4) constitute the CDCs. As we can begin to see more clearly, EnDQT has the benefit of providing a no non-local, retrocausal, or superdeterministic hidden-variable criterion for systems to have determinate values without modifying the fundamental equations of QT or appealing to relationalism.
In principle, the fundamental decoherence models will be grounded in interactions between the fundamental systems that have the DC. [^36] I am setting aside here exotic views that consider that such systems don’t exist. Since we are working in a non-relativistic regime, I have not provided such a model here, and I will be neutral about it. Also, given that we are also working in this regime, I haven’t provided a detailed relation between these different levels.
In Appendix A, I will also briefly provide a toy model to illustrate how EnDQT represents physical interactions. Readers are encouraged to read for more thorough details. That model will involve a series of interactions modeled via the spin-spin decoherence models.373737[109, 29]. These are the simplest decoherence models, and the example in the appendix is the simplest SDC model that one can conceive. I will also briefly argue that in a quantum field theoretic model, Lorentz and general covariance of the laws in principle will be respected. Furthermore, we will see that, in principle, any model of decoherence fulfilling CDC1)-CDC4) can be incorporated into this kind of SDC model. Here, in the main text, I will rather consider a simple and very schematic example. This example will involve systems $A$, $B$, and $C$, where $A$ is an initiator, in a toy mini-universe where the SDC that will be formed has the following structure, $A\rightarrow B\rightarrow C$. We assume that these will be quick interactions occurring at timescales where we can neglect the intrinsic evolution of the systems. For example, $A$ could be a system well approximated by a large set of quantum harmonic oscillators [^38] Or more precisely, bosonic modes. These interactions can be represented via the spin-boson mode. See, e.g., [58, 89]. or a set of spin-1/2 systems or two-level systems that can be approximated as spin-1/2 systems (see toy model in Appendix A). Each set of quantum systems interacts with a single spin-1/2 system. The collection of these systems constitutes system $B$. Then, system $C$ could be another two-level system that will interact with $B$ [^39] The interactions between the systems that constitute $B$ and system $C$ would be modeled via the so-called spin-spin decoherence models developed in [109, 29]. Let’s assume that $C$ starts interacting with the systems that constitute $B$ while $B$ are interacting with $A$ so that $B$ has the DC-$C$ (in agreement with the value mereology assumption), and $B$ can end up transmitting the DC to $C$ concerning some other system that $C$ might end up interacting with.
Moreover, in order for this interaction to fulfill CDC-ii) and to simplify, let’s assume that when $C$ starts interacting with $B$, it doesn’t change significantly the initial state that $B$ had when it started interacting with $A$. So, when $B$ and $C$ begin interacting, let’s neglect the evolution of the quantum states of $B$ while $A$ and $B$ interact, such that we can idealize that $B$ and $C$ start interacting only when the interaction between $A$ and $B$ ends. [^40] Another way to pose this assumption is to consider that while $A$ interacts with $B$, $C$ starts interacting with B in such a way that it doesn’t drive the states of $B$ out of being states that $A$ decoheres in the following sense: the Hamiltonian of interaction of $A$ and $B$ would still at least approximately commute with the (pointer) observable that these states are eigenstates of. Thus, we can just analyze the evolution of the quantum states of $A$ while $A$ and $B$ are interacting, where this interaction ends approximately at $t^{\prime}$ (in a certain relativistic reference frame). Let’s put a subscript SDC on the quantum states of a system if that system is an initiator or has the DC relative to some system belonging to an SDC. We then have the following interaction between $A$ and $B$,
@@
If $\left\langle E_{0}\left(t^{\prime}\right)\mid E_{1}\left(t^{\prime}\right)\right\rangle_{\text{A SDC}}\approx 0$ and $\left\langle E_{1}\left(t^{\prime}\right)\mid E_{0}\left(t^{\prime}\right)\right\rangle_{\text{A SDC}}\approx 0$ quasi-irreversibly when $A$ and $B$ end their interaction, we infer that $B$ has a determinate value of the (pointer) observable $monitored$ by $A$ at $t^{\prime}$ that arises from their interaction (i.e., let’s assume that is either 0 or 1) and acquires the DC-C. Let’s assume that $B$ has a determinate value 0. Now, let’s consider the interaction between $B$ and $C$, and assume that, given their interaction Hamiltonian, it ends at $t^{\prime\prime}$,
@@
The evolution of the interaction between $B$ and $C$ could be analyzed via the reduced density operator $\rho_{C}(t)$. Their interaction will allow $C$ to have a determinate value ($\uparrow$ or $\downarrow$) at $t^{\prime\prime}$ if $\langle E_{0}^{\prime\uparrow}\left(t^{\prime\prime}\right)|E_{0}^{\downarrow}\left(t^{\prime\prime}\right)\rangle_{B}\approx 0$ and $\langle E_{0}^{\prime\downarrow}\left(t^{\prime\prime}\right)|E_{0}^{\uparrow}\left(t^{\prime\prime}\right)\rangle_{B}\approx 0$ quasi-irreversibly when $B$ and $C$ end their interaction. $B$ will have a determinate value at $t^{\prime\prime}$ that arises from its interaction with $C$ where the possible values that it can have are represented via the eigenvalues of the observable that $|E_{0}^{\prime\uparrow}\left(t^{\prime\prime}\right)\rangle_{B}$ and $|E_{0}^{\prime\downarrow}\left(t^{\prime\prime}\right)\rangle_{B}$ are eigenstates of. Furthermore, $C$ can have the DC concerning some other system $D$ if it interacts with it before the interaction with $B$ ends. Note that since system $A$ is an initiator, if it is of the kind a), it has the DC concerning any system. So, if it is of this kind, its ability to give rise to other systems having determinate values and allowing them to have the DC doesn’t depend on its interactions with other systems.
@@
Figure 1: Directed graph that represents the transmission of the DC between systems.
@@
Figure 2: Directed graph involving common effects like the previous one, but just represents the whole system $S_{2}$.
Now, given the CDCs we can see how EnDQT justifies the assumptions 1)1)1 ) and 2)2)2 ) of decoherence models. Regarding 1)1)1 ), it is important that the Hamiltonian of interaction assumes a specific form to account for decoherence (given by the above commutativity criterion), which depends on the observables of the environment and the system because we can regard it as representing the law-like dynamics of SDCs, which leads to determinate values (more on this below when we assume a hypothesis that connects the pragmatic irreversible decoherence models with the fundamental decoherence models).
Regarding 2)2)2 ), the initial states of the systems (and associated coupling constants) are randomly distributed (which gives rise to the overlap terms going quasi-irreversibly to zero in decoherence models) because the subsystems of $E$ also need to belong to SDCs that indeterministically gave rise to the subsystems of $E$ having determinate values before interacting with $S$ and, hence, to a random distribution of their quantum states (and associated coupling constants). What about the distribution of the quantum states of the initiators? There are various possibilities, which will depend on what we regard as initiators, which are speculative. For example, that can be regarded as a brute fact of “the initial conditions of the universe” where initiators could be located. Or, given enough time, the quantum states of the initiators can achieve this overlap that triggers the process of propagation of the DC (more on this below). [^41] See also Appendix A.
As we will see, the explanation for why systems with the DC give rise to determinate values in this law-like way can be elegantly (as we will see) traced back to the initiators. So, one more benefit of adopting EnDQT with its CDCs is that it addresses some ad hoc or, in my view (at least), strange features of decoherence by seeing it as a tool to represent the dynamics of SDCs. In Appendix D, I go over a simple example of how EnDQT can model interference phenomena.
2.2 Two hypotheses
@@
I have been assuming that we can rely on fundamental decoherence to infer and represent how determinate values arise via SDCs. Fundamental decoherence represents the interactions between systems, starting with the initiators. The example above and CDC1)-CDC4) involve these models. So, fundamental decoherence represents interactions that render systems with determinate values in a certain situation involving an environment with a DC, given an appropriate and local Hamiltonian. Importantly, in order for a system $S$ to have a DC and lead other systems to have determinate values, it is plausible that it will typically have to belong to an environment with many systems that propagate the DC, which seems to lead to a process that is hard to reverse/control due to these many systems that have determinate values indeterministically. Furthermore, the more macroscopic is $S$, the harder this seems to be. I also explained that fundamental decoherence models seem to represent phenomena that resemble those represented by pragmatic irreversible decoherence models. Remember that pragmatic irreversible decoherence models are models that involve situations where it’s impossible to reverse the process they represent because they concern open/inaccessible environments or environments with many systems (but there is no reference to systems having the DC). It seems that we can use pragmatic irreversible decoherence models to infer when the processes represented by fundamental decoherence occur where the environmental systems have the DC. But, as we will see, more needs to be said about the behavior of SDCs and fundamental decoherence to ground the typically used pragmatic irreversible decoherence models as a proper inferential tool. More concretely, as an inferential tool to infer when systems have determinate values due to systems $E$ that have the DC in situations where we have not followed the interactions between systems that lead $E$ to have the DC.
@@
As a reminder, I have called the models of decoherence that don’t necessarily track the interactions involving systems with the DC pragmatic decoherence models. Another kind of pragmatic decoherence model is what I will call the pragmatic reversible decoherence models. These are models that represent a process that apparently involves decoherence in the sense that the overlap terms of the environment go quasi-irreversibly to zero. However, someone in some privileged position could reverse this process via operations on the system and environment, which is sometimes called recoherence. Or, to put it less pragmatically, they don’t involve enough degrees of freedom to be considered irreversible. This reversible process often occurs inside isolated environments or situations where the environmental degrees of freedom don’t become inaccessible to be reversed due to their in-practice isolation. Thus, the processes represented by these models aren’t what we typically consider to be decoherence.
@@
If we aren’t careful, the distinction between a pragmatic reversible model and an irreversible one may be ambiguous in some situations. The Wigner’s friend scenario424242[100]. is an example of a situation. Suppose an isolated lab occupies an arbitrarily large spatiotemporal region with a human agent inside (a “friend”). The lab is isolated in such a way that Wigner outside the lab can unitarily manipulate the state of the friend plus their target system that the friend interacts with, treating both as being in an entangled state. So, in this case, we could have an enormous lab with many systems getting entangled with the friend and their target system S for a long time. However, if we were Wigner, we wouldn’t consider that there was decoherence of S by the friend because Wigner could still unitarily reverse the state of the friend and their system. He would rather consider treating the friend plus target system interaction via a reversible decoherence pragmatic model. However, if the lab were open, he would treat their interaction via an irreversible decoherence pragmatic model. So, how do we exactly distinguish a reversible decoherence process from an irreversible one, since the reversible could also involve many systems? Let’s call this problem the Wigner’s friend ambiguity problem. As the discussion in the previous paragraphs suggests, if we want to address the measurement problem without changing the fundamental equations of QT, adding hidden variables to it, or adopting a relationalist view (more on this below), it seems that a way to deal with this problem is via paying attention to what constitutes an open environment and how it relates with SDCs.
@@
I have pointed out how fundamental decoherence involving members of an SDC seems to lead to processes represented by the pragmatic irreversible decoherence models. The first hypothesis aims to deal with the above ambiguity problem and ground the success of the pragmatic irreversible decoherence models in helping account for determinate values arising via SDCs in open environment situations, as opposed to the reversible models. These pragmatic irreversible decoherence models need to obey two plausible further constraints (given the local features of EnDQT): they need to represent local interactions (i.e., local), and they need to be empirically successful. Thus, I will hypothesize that
@@
The SDCs in our world are widespread in such a way that the pragmatic empirically successful local irreversible decoherence models in open environments track the interactions between systems that belong to SDCs serving as an environment for a target system that doesn’t belong to an SDC, but end up belonging to it. However, the SDCs in our world are also such that there can also exist processes represented via local empirically successful reversible decoherence pragmatic models, where the latter are tracking the interactions between systems that don’t belong to SDCs (SDCs-decoherence hypothesis).
@@
In other words, the SDCs-decoherence hypothesis states that the SDCs in our world are such that open environments will give rise to enough interactions between systems that have the DC and ones that may end up having it, allowing us to reliably approximate fundamental decoherence models using pragmatic irreversible decoherence models that concern open environments. From now on, I will just assume that processes represented by pragmatic irreversible decoherence models occur in open environments.
Under the plausible assumption that irreversible pragmatic decoherence models represent how the classical domain arises from quantum systems and that this classical domain is widespread, notice that the SDCs-decoherence hypothesis hypothesizes that systems tend to develop interactions that obey CDC1)-CDC4).
@@
On the other hand, it also hypothesizes that SDCs in our world are such that in isolated/controlled situations, we are able to isolate systems $E$ (which serve as “environments” for target systems in pragmatic decoherence models) from interacting with members of SDCs. They only interact with systems that don’t have the DC. $E$ will then interact with another system $S$, and this interaction is modeled via reversible pragmatic decoherence models.
So, given this hypothesis, if a pragmatic reversible decoherence model models specific situations $SI$ with predictive success, it allows us to infer the situations $SI$ where a system $S$ is interacting with environmental systems $E$ that don’t have the DC. Thus, we infer that in situations $SI$, no process that gives rise to determinate values occurs.